Solve using the qudratic equation.
w+4sqrt(w)-12=0

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Let . Then

Substitute:



Factor:





But means the positive square root, hence it cannot be negative, hence -6 is extraneous. Exclude.



Check:





Checks.


John


Answer by stanbon(75887)

Solve using the qudratic equation.
w+4sqrt(w)-12=0
---
Let w = k^2
---
k^2 + 4k -12 = 0
---
Factor:
(k+6)(k-2) = 0
k = -6 or k = 2
------------------------
Switching back to "w":
sqrt(w) = -6 or sqrt(w) = 2
w = 36 or w = 4
=============================
Checking these "answers" in w + 4sqrt(w) - 12 = 0.
w = 36 ? : 36 + 4*6 -12 = 0 ------------ false
w = 4 ? ; 4 + 4sqrt(4) - 12 = 0------------true
===============================
Final Answer: w = 4
=========================================
Cheers,
Stan H.